Allie has three fewer than twice the number of coins that Jonathan has. If Jonathan gave 2 coins to Allie, she would have three time as many coins as he would. How many coins does Allie have?
a. 2
b. 3
c. 5
d. 7
e. 9

Answer by ankor@dixie-net.com(15652)

Allie has three fewer than twice the number of coins that Jonathan has.
If Jonathan gave 2 coins to Allie, she would have three time as many coins as he would.
How many coins does Allie have?
:
Let a = no. of coins that A has
Let j = no. of coins that J has
:
Write an equation for each statement:
:
"Allie has three fewer than twice the number of coins that Jonathan has."
a = 2j - 3
:
"If Jonathan gave 2 coins to Allie, she would have three time as many coins as he would."
3(j-2) = a + 2
3j - 6 = a + 2
3j = a + 2 + 6
3j = a + 8
replace a with (2j-3)
3j = (2j-3) + 8
3j - 2j = -3 + 8
j = 5 coins has J originally
then
a = 2(5) - 3
a = 7 coins has Allie